OpenStudy (anonymous):

Find the area of a parallelogram with sides of 12 inches and 8 inches if one of the angles is 120 degrees. ____ times the square root of ___ sq. in.

OpenStudy (amistre64):

8*sin(120)*12 i think

OpenStudy (amistre64):

8*12 = 96 96 times sin(120)

OpenStudy (amistre64):

sin(120) = sin(60) = sqrt(3)/2 right?

OpenStudy (amistre64):

48 times sqrt(3) ?

OpenStudy (anonymous):

Correcto Mundo! Thanks! :)